# Cutoff for Contingency Table Random Walks

### v1

@@ 4.3 Strategy for Mixing Time Upper Bound
To form the mixing time upper bound, we will first show that the sum in (14) over $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$ is small enough. For this to be true, we need to assume that $q\gg n$. Specifically, we show that for $y$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$ that $\Phi_{Y}\left(\dfrac{2\pi y}{q}\right)\leq 1-\dfrac{1}{25nq}$. For the proof, we set the time to be
@@ 4.4 Bounding Decay in Characteristic Function for Large $\theta$
Next, note that $\Gamma$ has eigenvalue $1+\dfrac{1}{n-1}$ with eigenspace $\text{span}(\{\mathbbm{1}_{n-1}\})^{\perp}$ and multiplicity $n-2$, and eigenvalue $\dfrac{1}{n-1}$ with multiplicity $1$ and eigenvector $\mathbbm{1}_{n-1}$. Here, $\mathbbm{1}_{n-1}$ is the all-$1$ vector of dimension $n-1$. Let $\overline{\theta}:=\sum_{j=1}^{n-1}\theta_{j}/(n-1)$ be the average value of the coordinates of $\theta$. Suppose $\dfrac{1}{\sqrt{q}}<\|\theta\|_{\infty}\leq\dfrac{\pi}{2}$. If $|\overline{\theta}|\leq\dfrac{1}{2\sqrt{q}}$, then
@@ 4.5.1 Decomposing the Characteristic Function in $\ell^{2}$-Bound
as all the coordinates of $Y=(Y(1),\dots,Y(n-1))^{T}$ are zero except at most two coordinates which take values ±$\pm 1$. Combining (32) and (33), we obtain
@@ 4.5.3 Bounding the Sum for Small $y$
:=${\mathcal{M}}:=\left\{y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}:\left\|\dfrac{2\pi y}{q}\right\|_{\infty}\leq\dfrac{1}{\sqrt{q}}\right\}$
@@
$\leq\sum_{y\in{\mathcal{M}}}\exp\left(-\log(\alpha n)\left(\frac{1}{2}y_{1}^{2}+\sum_{k=2}^{n-1}(y_{k}-\mu_{k}(y))^{2}\frac{k}{k+1}\right)\right)$
$=\sum_{j_{1}=-N}^{N}\exp\left(-\frac{1}{2}\log(\alpha n)j_{1}^{2}\right)\left(\sum_{j_{2}=-N}^{N}\exp\left(-\frac{2}{3}\log(\alpha n)(j_{2}-\mu_{2}(y))^{2}\right)\right)\cdots$
Next, note that for all $12\leq k\leq n-$ 1, $\sum_{j=-N}^{N}\exp\left(-\dfrac{k+1}{k}\log(\alpha n)(j_{k}-x)^{2}\right)$ is maximized at $x=0$ as long as $n$ is large enough by Lemma 4. Thus we obtain
$\sum_{y\in{\mathcal{M}}}\exp(-\log(\alpha n)y^{T}\Gamma y)$ $\leq\left(\sum_{j=-N}^{N}\exp\left(-\frac{1}{2}\log(\alpha n)j^{2}\right)\right)\prod_{k=2}^{n-1}\left(\sum_{j=-N}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)j^{2}\right)\right)$ (42)
@@
We bound the inner expression as follows:
$\sum_{j=-N}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)j^{2}\right)$ (43)
$=1+2\sum_{j=1}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)j^{2}\right)$
[…]
@@
Using this bound, the fact that $\log(1+x)\leq x$ for all $x>-1$, and Lemma 2, we have
@@ 4.6 Mixing Time Upper Bound
Putting together (14), (31), and (44), we conclude that when $t=\dfrac{nq^{2}\log(\alpha n)}{8\pi^{2}}\left(1-\dfrac{1}{3q}\right)^{-1}$,
@@ 4.7 Mixing Time Lower Bound
We will now compute the mixing time lower bound for the $1\times n$ contingency table walk in continuous time. We showed that the discrete-time increment $Y$ has equivariant coordinates, with each coordinate having a marginal variance of $\dfrac{2}{n}$. Recall that the off-diagonal entries of the correlation matrix are equal to exactly $\dfrac{-1}{n-1}$. Since at each jump the $1\times n$ contingency table random walk updates at most two coordinates, each by ±$\pm 1$, we have that $\|Y\|_{1}\leq 2$. And therefore, applying (13) and setting $r=2$, $\sigma^{2}=\dfrac{2(n-1)}{n}$, we obtain
@@ 5.1 Setup
Let $\{Y_{m}\}_{m\in{\mathbb{N}}}$ be an arbitrary irreducible symmetric discrete-time Markov chain on the torus ${\mathbb{Z}}_{q}^{n}$ with transition matrix $P$, such that
${\mathbb{P}}(Y_{m}=y_{m}|Y_{m-1}=y_{m-1})=\mu(y_{m}-y_{m-1}).$
Suppose that $\|Y\|_{1}\leq r$, $\|Y\|_{\infty}<q/2$, and $r^{2}\ll q$. We set $Y_{0}=0$. Let $Y\in{\mathbb{Z}}^{n}$ be a lattice-valued random vector with law $\mu$ mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$. Note that $Y$ is symmetric. Assume $Y$ has the same marginal variance $\dfrac{\sigma^{2}}{n}$ for each coordinate. Write $\operatorname{Var}(Y)=\dfrac{\sigma^{2}}{n}\Gamma$ where $\Gamma$ is the correlation matrix of $Y$. Let $\{X_{t}\}$ be the rate-1 continuous-time Markov chain on ${\mathbb{Z}}_{q}^{n}$ with rate matrix $Q=P-I$ and transition kernel $H_{t}=\exp(tQ)$.
@@
$\left\{\exp\left(t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}\right\}.$ (45)
@@
Using the same strategy in Section 4, we will upper bound the sum above in two regimes of $y$ separately.
@@
When $y$ is big, we need $Y$ to satisfy the following decay condition. Note that for this condition to hold, we may need some extra asymptotic assumptions on $q$.
@@
:=${\mathcal{L}}:=\left\{y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}:\left\|\dfrac{2\pi y}{q}\right\|_{\infty}>\dfrac{1}{\sqrt{q}}\right\}.$
@@
We give a few examples in which the Decay Condition is met.
@@
If $\{Y_{m}\}$ is the simple random walk on ${\mathbb{Z}}_{q}^{n}$, then $Y$ is the increment for the simple random walk on ${\mathbb{Z}}^{n}$. In other words, $Y$ takes values in $\{(0,\dots,0,\pm 1,0\dots,0)\in{\mathbb{Z}}^{n}\}$, each with probability $\dfrac{1}{2n}$. In this case, $r=\sigma^{2}=1$. The characteristic function $\Phi_{Y}(\theta)=\dfrac{1}{n}\sum_{k=1}^{n}\cos(\theta_{k})$ for $\theta=(\theta_{1},\dots,\theta_{n})\in{\mathbb{R}}^{n}$. Using the bound (18) for cosine, we have that for $y=(y_{1},\dots,y_{n})\in{\mathcal{L}}$ and $t\geq\dfrac{nq^{2}\log(n)}{4\pi^{2}}$,
@@
$\leq\exp\left(2t\left(-\frac{2}{25n}\frac{4\pi^{2}y_{k}^{2}}{q^{2}}\right)\right)$
$\leq\exp\left(2t\left(-\frac{2}{25nq}\right)\right)$
$\leq n^{-\dfrac{q}{25\pi^{2}}}.$
@@
Let $f$ be a real valued function function on a state space $\Omega$, and let $P$ be a reversible and irreducible transition matrix on $\Omega$ with stationary distribution $\pi$. We define the Dirichlet form,
$\mathcal{E}(f)=\langle(I-P)f,f\rangle_{\pi}$
where we define the inner product,
$\langle f,g\rangle_{\pi}=\sum_{x\in\Omega}\pi(x)f(x)g(x)$
For any non-constant eigenfunction $f$ of $P$ with $\operatorname{Var}_{\pi}(f)=1$, we have that,
@@
Using this, let $P$ be the transition matrix for $\{Y_{m}\}$ and $\widetilde{P}$ be the transition matrix for the simple random walk on ${\mathbb{Z}}_{q}^{n}$. Suppose we have a constant $B=B(n)$ so that for any real valued function $g$ on $\Omega$ we have,
~$\widetilde{\mathcal{E}}(g)\leq B\mathcal{E}(g)$ (48)
@@
Let $\{Y_{m}\}$ again be an irreducible symmetric random walk on ${\mathbb{Z}}_{q}^{n}$ as in the setup above in Section 5.1, and let $Y$ be defined accordingly. Let $\{S_{m}\}$ be a random walk on ${\mathbb{Z}}^{n}$ with increment $Y$. First, we observe that (45) with $t=1$ corresponds to the characteristic function for a random variable on $\mathbb{Z}^{n}$ with probability mass function $p$ given by
$p(x)=e^{-1}\sum_{m=0}^{\infty}\dfrac{\mathbb{P}(S_{m}=x)}{m!}.$
By the irreducibility of $S_{m}$, we know that the random walk on $\mathbb{Z}^{n}$ with incremental distribution $p$ is aperiodic and irreducible. Therefore, by Lemma 2.3.2 of [16], for all $\theta\in[-\pi,\pi]^{n}$,
@@
On the other hand, when $y$ is small, we will assume the following correlation condition on the correlation matrix $\Gamma$ of $Y$.
@@
    where $\Psi_{n,k}^{(11)}$ is submatrix attained by taking the first $k-1$ rows and columns of $\Psi_{n,k}^{(11)}$, and $\Psi_{n,k}^{(22)}$ is the $(k,k)$ entry of $\Psi_{n,k}$. Let $a_{1}^{(n)}:=\Psi_{n,1}^{-1}$, and for $k\geq 2$, let
@@
:=${\mathcal{M}}:=\left\{y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}:\left\|\dfrac{2\pi y}{q}\right\|_{\infty}\leq\dfrac{1}{\sqrt{q}}\right\}.$
If the correlation matrix $\Gamma$ of $\xi_{1}$ satisfies the Correlation Condition with bounding function $g(\alpha)$, then
@@
5.4 Cutoff for a Family of Random Walks on ${\mathbb{Z}}_{q}$
We now summarize our generalization in Sections 3 and 5 with the following theorem.

### v2

@@ 4.3 Strategy for Mixing Time Upper Bound
We will form the mixing time upper bound via the $\ell^{2}$ bound in (18). We use the link between the eigenvalues of $X_{t}$ and the characteristic function of the lifted distribution to gain control on the rate of decay of the spectrum of $X_{t}$. This is done by Taylor expanding $\Phi_{Y}(2\pi y/q)$ in the regime where $\rVert y\rVert_{\infty}\leq\dfrac{\sqrt{q}}{2\pi}$. In this regime, the eigenvalues of $X_{t}$ resemble the Gaussian characteristic function and depend on the quadratic form $y^{T}\Gamma y$, where $\Gamma$ is the correlation matrix of the lifted distribution $Y$. In Section 4.5.2., we simplify the quadratic form $y^{T}\Gamma y$ based on Schur complements of $\Gamma^{-1}$. We then use this in Section 4.5.3. to show the $\ell^{2}$-bound is sufficiently small in the high frequency regime ($\rVert y\rVert_{\infty}\leq\dfrac{\sqrt{q}}{2\pi}$). In Section 4.5.1 we also show all contributions to the $\ell^{2}$-bound in the lower frequency regime ($\rVert y\rVert_{\infty}>\dfrac{\sqrt{q}}{2\pi}$) by directly analyzing the decay of $\Phi_{Y}$, the characteristic function of our lifted distribution, outside of a neighborhood of zero under the regime where $q\gg n$.
We will first show that the sum in (18) over $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$ is small enough. For this to be true, we need to assume that $q\gg n$. Specifically, we show that for $y$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$ that $\Phi_{Y}\left(\dfrac{2\pi y}{q}\right)\leq 1-\dfrac{1}{25nq}$. For the proof, we set the time to be
@@ 4.4 Bounding Decay in Characteristic Function for Large $\theta$
Next, note that $\Gamma$ has eigenvalue $1+\dfrac{1}{n-1}$ with eigenspace $\text{span}(\{\mathbbm{1}_{n-1}\})^{\perp}$ and multiplicity n2n-2, and eigenvalue $\dfrac{1}{n-1}$ with multiplicity $1$ and eigenvector $\mathbbm{1}_{n-1}$. Here, $\mathbbm{1}_{n-1}$ is the all-$1$ vector of dimension n1n-1. Let $\overline{\theta}:=\sum_{j=1}^{n-1}\theta_{j}/(n-1)$ be the average value of the coordinates of $\theta$. Suppose $\dfrac{1}{\sqrt{q}}<\|\theta\|_{\infty}\leq\dfrac{\pi}{2}$. If $|\overline{\theta}|\leq\dfrac{1}{2\sqrt{q}}$, then
@@ 4.5.1 Decomposing the Characteristic Function in $\ell^{2}$-Bound
as all the coordinates of $Y=(Y(1),\dots,Y(n-1))^{T}$ are zero except at most two coordinates which take values $\pm 1$. Combining (36) and (37), we obtain
@@ 4.5.3 Bounding the Sum for Small $y$
${\mathcal{M}}:=\left\{y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}:\left\|\dfrac{2\pi y}{q}\right\|_{\infty}\leq\dfrac{1}{\sqrt{q}}\right\}$
@@
$\sum_{y\in{\mathcal{M}}}\exp(-\log(\alpha n)y^{T}\Gamma y)$
$\leq\sum_{y\in{\mathcal{M}}}\exp\left(-\log(\alpha n)\left(\frac{1}{2}y_{1}^{2}+\sum_{k=2}^{n-1}(y_{k}-\mu_{k}(y))^{2}\frac{k}{k+1}\right)\right).$
By regrouping the summation with the definition of $\mathcal{M}$, the expression above can be written as
$\sum_{j_{1}=-N}^{N}\exp\left(-\frac{1}{2}\log(\alpha n)j_{1}^{2}\right)\left(\sum_{j_{2}=-N}^{N}\exp\left(-\frac{2}{3}\log(\alpha n)(j_{2}-\mu_{2}(y))^{2}\right)\right)\cdots$ (45)
Next, note that for all $2\leq k\leq n-1$, $\sum_{j=-N}^{N}\exp\left(-\dfrac{k+1}{k}\log(\alpha n)(j_{k}-x)^{2}\right)$ is maximized at $x=0$ as long as $n$ is large enough by Lemma 4. Thus we obtain
@@
$\leq\left(\sum_{j=-N}^{N}\exp\left(-\frac{1}{2}\log(\alpha n)j^{2}\right)\right)\prod_{k=2}^{n-1}\left(\sum_{j=-N}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)j^{2}\right)\right)$
@@
We look at the inner expression:
$\sum_{j=-N}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)j^{2}\right)=1+2\sum_{j=1}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)j^{2}\right)$
Bounding the sum by an integral, we have that the previous expression is at most
$1+2\exp\left(-\frac{k}{k+1}\log(\alpha n)\right)+2\int_{1}^{N}\exp\left(-\frac{k}{k+1}\log(\alpha n)z^{2}\right)dz$ (47)
$\leq 1+2\exp\left(-\frac{k}{k+1}\log(\alpha n)\right)+2\int_{1}^{\infty}z\cdot\exp\left(-\frac{k}{k+1}\log(\alpha n)z^{2}\right)dz$
@@
Using this bound, the fact that $\log(1+x)\leq x$ for all x>1x>-1, and Lemma 2, we have
@@ 4.6 Mixing Time Upper Bound
Putting together (18), (35), and (48), we conclude when $t=\dfrac{nq^{2}\log(\alpha n)}{8\pi^{2}}\left(1-\dfrac{1}{3q}\right)^{-1}$,
@@ 4.7 Mixing Time Lower Bound
We will now compute the mixing time lower bound for the $1\times n$ contingency table walk in continuous time. We showed that the discrete-time increment $Y$ has equivariant coordinates, with each coordinate having a marginal variance of $\dfrac{2}{n}$. Recall that the off-diagonal entries of the correlation matrix are equal to exactly $\dfrac{-1}{n-1}$. Since at each jump the $1\times n$ contingency table random walk updates at most two coordinates, each by $\pm 1$, we have that $\|Y\|_{1}\leq 2$. And therefore, applying (17) and setting $r=2$, $\sigma^{2}=\dfrac{2(n-1)}{n}$, we obtain
@@ 5.1 Setup
Let $\{Y_{s}\}_{s\in{\mathbb{N}}}$ be an arbitrary irreducible symmetric discrete-time Markov chain on the torus ${\mathbb{Z}}_{q}^{n}$ with transition matrix $P$, such that
${\mathbb{P}}(Y_{s}=y_{s}|Y_{s-1}=y_{s-1})=\mu(y_{s}-y_{s-1}).$
We set $Y_{0}=0$. Let $Y\in{\mathbb{Z}}^{n}$ be a lattice-valued random vector with law $\mu$ mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$. Suppose that $\|Y\|_{1}\leq r$, $\|Y\|_{\infty}<q/2$, and $r^{2}\ll q$. Note that $Y$ is symmetric. Assume $Y$ has the same marginal variance $\dfrac{\sigma^{2}}{n}$ for each coordinate. Write $\operatorname{Var}(Y)=\dfrac{\sigma^{2}}{n}\Gamma$ where $\Gamma$ is the correlation matrix of $Y$. We will assume that $\Gamma$ is positive definite. Let $\{X_{t}\}$ be the rate-1 continuous-time Markov chain on ${\mathbb{Z}}_{q}^{n}$ with rate matrix Q=PIQ=P-I and transition kernel $H_{t}=\exp(tQ)$.
@@
$\left\{\exp\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}\right\}.$ (49)
@@
We will again upper bound the sum above in two regimes of $y$ separately.
Our strategy in generalizing our upper bound method will be similar to that in Section 4 by analyzing the $\ell^{2}$ bound. We leverage the correspondence between eigenvalues of our walk $X_{t}$ and the characteristic function of its lift, $\Phi_{Y}$. Similar to a local central limit theorem, we show our characteristic function $\Phi_{Y}$ is sufficiently decayed in the low frequency regime $\rVert y\rVert_{\infty}\geq\dfrac{q}{2\pi}$, and thus our $\ell^{2}$ sum is controlled. We define a Decay Condition that describes the amount of control needed, and briefly discuss tools to show the condition holds. In the high frequency regime $\rVert y\rVert_{\infty}\leq\dfrac{\sqrt{q}}{2\pi}$, we Taylor expand $\Phi_{Y}$ and leverage the correlation structure of the lifted distribution $Y$. We form a Correlation Condition based on $\Gamma$ to gain control on the rate of decay of the spectrum of $X_{t}$.
@@
When $y$ is big, we need $Y$ to satisfy the following decay condition, so that we can show that our $\ell_{2}$ sum is very small in the low frequency regime in a format analogous to a local central limit theorem. This condition is based on the rate of decay of the characteristic function $\Phi_{Y}$ outside of a neighborhood of zero. For this condition to hold, we may need some additional asymptotic assumptions on $q$.
@@
${\mathcal{L}}:=\left\{y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}:\left\|\dfrac{2\pi y}{q}\right\|_{\infty}>\dfrac{1}{\sqrt{q}}\right\}.$
@@
Note that we required $q\gg n$ in the $1\times n$ contingency table case for this condition to hold. We give a few other examples in which the Decay Condition is met.
@@
If $\{Y_{s}\}_{s\in{\mathbb{N}}}$ is the simple random walk on ${\mathbb{Z}}_{q}^{n}$, then $Y$ is the increment for the simple random walk on ${\mathbb{Z}}^{n}$. In other words, $Y$ takes values in
$\{(0,\dots,0,\pm 1,0\dots,0)\in{\mathbb{Z}}^{n}\},$
each with probability $\dfrac{1}{2n}$. In this case, $r=\sigma^{2}=1$. The characteristic function $\Phi_{Y}(\theta)=\dfrac{1}{n}\sum_{k=1}^{n}\cos(\theta_{k})$ for $\theta=(\theta_{1},\dots,\theta_{n})\in{\mathbb{R}}^{n}$. Using the bound (22) for cosine, we have that for $y=(y_{1},\dots,y_{n})\in{\mathcal{L}}$ and $t\geq\dfrac{nq^{2}\log(n)}{4\pi^{2}}$,
@@
$=\exp\left(-2t\left(\sum_{k=1}^{n}\frac{2}{25n}\frac{4\pi^{2}y_{k}^{2}}{q^{2}}\right)\right)$
For $y\in\mathcal{L}$, $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$. Since $\sum_{k=1}^{n}y_{k}^{2}\geq\|y\|_{\infty}^{2}$, the expression above can be upper bounded by
$\exp\left(2t\left(-\frac{2}{25n}\frac{4\pi^{2}\|y\|_{\infty}^{2}}{q^{2}}\right)\right)\leq\exp\left(2t\left(-\frac{2}{25nq}\right)\right)\leq n^{-\dfrac{q}{25\pi^{2}}}.$
@@
Let $f$ be a real-valued function on a state space $\Omega$, and let $P$ be a reversible and irreducible transition matrix on $\Omega$ with stationary distribution $\pi$. We define the Dirichlet form,
$\mathcal{E}(f):=\langle(I-P)f,f\rangle_{\pi}$
where the inner product is defined as
$\langle f,g\rangle_{\pi}:=\sum_{x\in\Omega}\pi(x)f(x)g(x)$
For any non-constant eigenfunction $f$ of $P$ with $\operatorname{Var}_{\pi}(f)=1$, we have that
@@
Using this, let $P$ be the transition matrix for $\{Y_{s}\}$ and $\widetilde{P}$ be the transition matrix for the simple random walk on ${\mathbb{Z}}_{q}^{n}$. Suppose we have a constant $B=B(n)$ so that for any real-valued function $g$ on $\Omega$ we have
$\widetilde{\mathcal{E}}(g)\leq B\mathcal{E}(g)$ (52)
@@
Let $\{Y_{s}\}_{s\in{\mathbb{N}}}$ again be an irreducible symmetric random walk on ${\mathbb{Z}}_{q}^{n}$ as in the setup above in Section 5.1, and let $Y$ be defined accordingly. Let $\{S_{s}\}_{s\in{\mathbb{N}}}$ be a random walk on ${\mathbb{Z}}^{n}$ with increment $Y$. First, we observe that (49) with $t=1$ corresponds to the characteristic function for a random variable on $\mathbb{Z}^{n}$ with probability mass function $p$ given by
$p(x)=e^{-1}\sum_{s=0}^{\infty}\dfrac{\mathbb{P}(S_{s}=x)}{s!}.$
By the irreducibility of $S_{s}$, we know that the random walk on $\mathbb{Z}^{n}$ with incremental distribution $p$ is aperiodic and irreducible. Therefore, by Lemma 2.3.2 of [17], for all $\theta\in[-\pi,\pi]^{n}$,
@@
On the other hand, when $y$ is small, we will assume the following correlation condition on the correlation matrix $\Gamma$ of $Y$. Our correlation condition depends on the magnitude of Schur complements of $\Psi$, the inverse of the random walk’s correlation matrix $\Gamma$. Schur complements of $\Psi$ have a direct connection to the conditional variances of the multivariate Gaussian distribution with correlation matrix $\Gamma$. This connection to the multivariate Gaussian distribution arises from studying the lifted random walk of $X_{t}$, which then is linked to the local limit theorem in our regime when $q$ is large. Our condition checks the magnitude of conditional variances for the multivariate Gaussian distribution with correlation matrix $\Gamma$. We will now state the condition and afterwards provide more interpretation.
@@
    where $\Psi_{n,k}^{(11)}$ is the submatrix attained by taking the first k1k-1 rows and columns of $\Psi_{n,k}$, and $\Psi_{n,k}^{(22)}$ is the $(k,k)$ entry of $\Psi_{n,k}$. Let $a_{1}^{(n)}:=\Psi_{n,1}^{-1}$, and for $k\geq 2$, let
@@
As previously mentioned, the Correlation Condition depends on conditional variances of the multivariate Gaussian distribution. First, note that for $1\leq i\leq n$, the $i$th diagonal entry of $\Psi$ can be written as $\operatorname{Var}(Z_{i}\mid Z_{1},\ldots,Z_{i-1},Z_{i+1},\ldots,Z_{n})^{-1}$, where $Z$ has the multivariate Gaussian distribution with correlation matrix $\Gamma$, assuming unit variances. Thus, the condition first checks that the variance of one coordinate conditioned on the rest stays bounded above from zero as the dimension $n$ grows to infinity. Moreover, the elements $a^{(n)}_{k}$ are equal to the conditional variances $\operatorname{Var}(Z_{k}\mid Z_{k+1},\dots Z_{n})$, where $Z$ again has the multivariate Gaussian distribution with correlation matrix $\Gamma$ and unit variances. The condition therefore checks that conditional variance of the multivariate Gaussian distribution with correlation matrix $\Gamma$ also decays from $1$ sufficiently slow as we condition on more and more variables.
@@
${\mathcal{M}}:=\left\{y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}:\left\|\dfrac{2\pi y}{q}\right\|_{\infty}\leq\dfrac{1}{\sqrt{q}}\right\}.$
If the correlation matrix $\Gamma$ of $\xi_{1}$ satisfies the Correlation Condition with the bounding function $g(\alpha)$, then
@@
5.4 Cutoff for a Family of Random Walks on ${\mathbb{Z}}_{q}^{n}$
We now summarize our generalization in Sections 3 and 5 with the complete statement of Theorem 3.